Hello, I need help with the following questions. I referred to numerous sources, but could not...

Hello,

I need help with the following questions. I referred to numerous sources, but could not calculate the solutions for these two questions. Please show step by step calculations. Thank you!

  1. Given a Correlation of Coefficient of 0.42, what is the ratio of SSR/SST?
    1. 0.00
    2. 0.42
    3. 0.48
    4. 0.18
  1. A regression equation is computed to be: Y = 15.8 + 11.2X.
    We know the corresponding SST = 685 and SSR = 282. What is the correlation coefficient?
    1. 0.83
    2. 0.41
    3. 0.64
    4. Cannot be computed from data available
Answer #1

Answer - 1: As we know:

Coefficient of determination (R2) = SSR/SST

Hence,

SSR/SST = (0.42)2 = 0.1764 = 0.18

Option D is correct.

Answer - 2: R2 = SSR/SST = 282/685 = 0.4117

Also, the sign of correlation is same as sign of slope. In this case, slope is +11.2 so correlation coefficient will also be positive.

Hence,

Correlation coefficient (R) = = 0.64

Option C is correct.